Jason bought x pens with a total amount of RM48.He gave away 3 pens to his brother as birthday present.He plans to make a profit of RM3 from each of the remaining pens by selling them to his friends.Jason managed to make a total profit of RM15from the sales.Based on the above information ,find the value of x.
1
Expert's answer
2021-05-11T11:38:24-0400

If Jason managed to make a total profit of RM15 from the sales, and the profit per unit is RM3, then he sold 15/3 = 5 units. He gave away 3 pens to his brother as birthday present, then the initial amount x = 5 + 3 = 8 units.